Three consecutive even integers are such that the square of the third is 100 more than the square of the second. Find the three integers.
Answer by checkley75(3666)

let x, (x+2) & (x+4) be the three consecutive even integers.
(x+4)^2=(x+2)^2+100
x^2+8x+16=x^2+4x+4+100
x^2-x^2+8x-4x=100+4-16
4x=88
x=88/4
x=22 ans. for the first integer.
22+2=24 for the middle integer.
22+4=26 for the third integer.
Proof:
26^2=24^2+100
676=576+100
676=676